# The Nature of Turbulence at Sub-Electron Scales in the Solar Wind

### v1

@@ Abstract
The nature of turbulence at sub-electron scales has remained an open question, central to understanding how electrons are heated in the solar wind. This is primarily because spacecraft measurements have been limited to magnetic field fluctuations alone. We resolve this by deriving new high-resolution density fluctuations from spacecraft potential measurements of Parker Solar Probe resolving scales smaller than the electron gyro-radius ($\rho_{e}$). A systematic comparison of the density and magnetic spectra shows that both steepen near the electron scales. Notably, the density spectrum exhibits slopes close to 10/3-10/3, while the magnetic spectrum becomes consistently steeper than the density spectrum at scales smaller than $\rho_{e}$, indicating that the turbulence becomes electrostatic. These results are consistent with theoretical predictions of an electron entropy cascade, which may explain the irreversible dissipation of turbulent energy at sub-$\rho_{e}$ scales.
@@ Introduction:
Turbulence in the solar wind drives a nonlinear cascade of energy injected at large scales to smaller kinetic scales, where it gets dissipated [1, 2, 3, 4]. This is well established from observations of magnetic power spectra, which exhibit a broadband wavenumber power-law scaling $P(k)\sim k^{\alpha}$ with $\alpha\simeq-3/2$ to 5/3-5/3 at magneto-hydrodynamic (MHD) scales [5, 6, 7, 8], consistent with models of Alfvénic turbulence [9, 10, 11, 12]. At sub-ion scales, between the ion gyro-radius ($\rho_{i}$) and electron inertial scale ($d_{e}$), density ($n$) and magnetic field ($B$) spectra have been extensively measured, showing slopes close to 7/3-7/3 or 8/3-8/3 [13, 14], in agreement with predictions of kinetic Alfvén (KA) turbulence [15, 16, 17, 18, 19, 20, 21]. In the range between $d_{e}$ and $\rho_{e}$, an increase in magnetic compressibility was found, with the magnetic spectrum showing a slope near 11/3-11/3, interpreted as a transition from KA to inertial-KA cascade under low electron beta ($\beta_{e}$) conditions [22]. In contrast, instrumental limitations have prevented a similar extent of progress at sub-electron scales (below the electron gyro-radius, $\rho_{e}$), with measurements affected by low signal-to-noise ratio (SNR) and restricted to magnetic fluctuations alone. As a result, the nature of turbulence at these scales has remained an open question.
@@
We use high-resolution search coil magnetometer (SCM) and antenna potential DFB burst mode data from the FIELDS instrument suite on PSP, sampled at $\sim 19\,\mathrm{kHz}$ [43, 44, 45]. Data from encounters E10–E18 were surveyed for intervals with enhanced fluctuation amplitudes near the Sun, where density and magnetic spectra could be resolved at scales below $\rho_{e}$. Although $\rho_{e}$ and $d_{e}$ shift beyond instrumental resolution near perihelion, several $27\,\mathrm{s}$ intervals within $0.15–0.2\,\textrm{au}$ were identified where sub-$\rho_{e}$ scales could be resolved. We selected 76 such intervals without dust hits or wave activity, exhibiting $\textrm{SNR}\gtrsim 10$ at the observed spectral break around $d_{e}$ (detailed in the next section). The intervals have mean plasma parameters: $\beta_{i}=\beta_{e}\simeq 0.4$ and temperature ratio, $T_{i}/T_{e}\simeq 1$.
Since SPAN, SPC [46], and QTN [47] density measurements have a low sampling rate ($\sim 0.14\,\mathrm{Hz}$), we use spacecraft potential ($V_{\textrm{sc}}$) as a proxy for density to resolve scales smaller than $\rho_{e}$. This method is based on the following principle: Solar illumination charges the spacecraft body positive via photoemission, attracting nearby electrons and thereby making $V_{\mathrm{sc}}$ sensitive to local electron density ($n$) [48, 49, 50, 51]. The resulting balance between photoelectric and thermal electron currents leads to an exponential relation: $n\propto e^{-V_{\mathrm{sc}}/V_{\mathrm{pe}}+C_{\mathrm{b}}}$, where $V_{\mathrm{pe}}$ is the photoelectron thermal energy. $V_{\mathrm{sc}}$ is computed as the negative average of the four PSP antenna potentials, $V_{\mathrm{sc}}=-(V_{1}+V_{2}+V_{3}+V_{4})/4$, which includes the antenna bias voltage. To account for this bias, an offset parameter $(C_{\mathrm{b}})$ was included. Other currents, such as ion or secondary electron currents, may be present but are negligible. This relation is indeed observed between electron QTN density measurements and $V_{\mathrm{sc}}$ for the intervals used. Typically at sub-ion scales, $\delta V_{\textrm{sc}}/V_{\mathrm{pe}}\ll 1$, implying $\delta n\propto\delta V_{\mathrm{sc}}$, up to first order. Thus, the two spectra share the same shape and the $V_{\textrm{sc}}$ spectrum can be used to measure the density spectral slopes.
@@
For the selected intervals, the power spectral density (PSD) of $B$ and $V_{\textrm{sc}}$ were computed using fast Fourier transforms. Fig. 1(a,b) shows the $V_{\textrm{sc}}$ and trace $B$ spectra vs spacecraft frame frequency ($f$) for a typical interval on 2022-06-04, 20:54:10–20:54:37. The trace $B$ spectrum is computed using only the $y$ and $z$ SCM components as the $x$ component is unavailable. For this interval, the low-frequency $B$ spectrum using the flux-gate magnetometer (MAG) data is also shown (MAG data was rotated into SCM coordinates before computing the trace [53]). The spectra represented are smoothed using a variable-width running mean window. Fig. 1(c) shows the $V_{\textrm{sc}}$ and $B$ spectra compensated by the sub-ion range scaling. The Doppler shifted frequencies corresponding to $k_{\perp}d_{e}=1$ and $k_{\perp}\rho_{e}=1$ are marked using, $f_{d_{e}}(f_{\rho_{e}})\simeq V_{\mathrm{rl}}\sin\theta_{BV_{\mathrm{rl}}}/2\pi d_{e}(\rho_{e})$, assuming Taylor’s hypothesis [54]. Here, $V_{\mathrm{rl}}$ is the magnitude of the relative velocity between the solar wind and the spacecraft, and $\theta_{BV_{\mathrm{rl}}}$ is the angle between $\mathbf{B}$ and $\mathbf{V_{\mathrm{rl}}}$. A piecewise power-law,
$P(f)=\begin{cases}A_{1}f^{\alpha_{1}},&f<f_{b}\\ A_{2}f^{\alpha_{2}},&f\geq f_{b}\end{cases}$ (1)
@@
Figure 1: Power spectral density of (a) $V_{\mathrm{sc}}$ and (b) $B$ (MAG - dashed, SCM - solid). The least square fits and the break frequency ($f_{b}$) using Eq. 1 are shown. Thin-dashed lines indicate the flat DFB and the frequency dependent SCM noise floors. (c) $V_{\mathrm{sc}}$ and $B$ spectra compensated by the sub-ion range scaling $(f^{-\alpha_{1}}P(f))$. The compensated spectra are shown as dotted lines for SNR $\lesssim 2$.
@@
Figure 2: (a) $V_{\mathrm{sc}}$ and (b) $B$ spectra for all intervals. Average slopes at sub-ion and sub-electron scales are shown. Blue diamonds mark the breaks obtained from fits. Mean $f_{d_{e}}$ and its variation across the intervals are shown by the red marker.
As depicted in Fig. 1, the $n$ and $B$ spectra show similar slopes of 2.77-2.77 and 2.80-2.80, respectively, in the sub-ion range ($f<f_{d_{e}}$). A spectral break is identified from the fit at frequency $f_{b}$ approximately near $f_{d_{e}}$ or $f_{\rho_{e}}$ in both spectra. Thereafter, the $B$ spectrum steepens to 4.40-4.40 while the $n$ spectrum is comparatively shallower at 3.33-3.33. This transition is clear from the compensated spectra, which steepen compared to the sub-ion range scaling ($f^{\alpha_{1}}$, horizontal dashed line) at $f_{d_{e}}$ or $f_{\rho_{e}}$. A similar trend was observed across all studied intervals as evident from Fig. 2. This figure also marks the spectral breaks (blue diamonds), showing they appear approximately around $f_{d_{e}}$. While a steepening in the $B$ spectrum has been reported before close to the electron scales, we show here that the $n$ spectrum also steepens, although to a lesser degree than the $B$ spectrum. This difference is evident from the distribution of $n$ (Fig. 3(a)) and $B$ (Fig. 3(b)) spectral slopes obtained from the fits (Eq. 1) across all intervals. In the sub-ion range ($f<f_{b}$), the distributions peak near 8/3-8/3, with mean slopes of $2.59\pm 0.02$ for $B$ and $2.55\pm 0.03$ for $n$, with the uncertainties indicating the standard error of the mean. This is consistent with previous observations [13, 14], and the theoretical and numerical predictions of KA turbulence, which suggest a slope close to 7/3-7/3 or 8/3-8/3 for both fields [15, 16, 17, 18, 19, 20, 21]. In the sub-electron range ($f>f_{b}$), both spectra steepen, but differ in slope. While $n$ shows a mean slope of $3.20\pm 0.03$, $B$ steepens to $3.78\pm 0.04$, occasionally reaching up to 4.8-4.8. This is the main result of this Letter, with this difference in slope indicating a breakdown of KA turbulence close to $\rho_{e}$.
@@ Conclusion:
We have presented high-resolution density spectra obtained from spacecraft potential measurements, alongside magnetic spectra resolving scales smaller than $\rho_{e}$, leading to the following main findings: At sub-$\rho_{e}$ scales, the density and magnetic spectra diverge, with the density spectrum exhibiting slopes close to 10/3-10/3 and the magnetic spectrum comparatively steeper. These results indicate that turbulence at these scales becomes predominantly electrostatic and represents key evidence for the presence of an electron entropy cascade [37, 20], a phase-space process, which can explain how electrons are irreversibly heated in a wide range of space and astrophysical systems. In this context, our study also motivates the need for improved particle instruments to directly measure the associated small-scale velocity-space structures and to fully understand the role of nonlinear phase-space dynamics in the irreversible dissipation of turbulent energy in weakly collisional plasmas.
@@ Acknowledgements.
S.M. is supported by a QMUL PhD studentship. C.H.K.C. is supported by UKRI Future Leaders Fellowship MR/W007657/1. C.H.K.C. and D.M. are supported by STFC Consolidated Grant ST/X000974/1. SM acknowledges A.A. Schekochihin, P.A. Simon, C. Sishtla, S. Greess and F. Koller for useful discussions. Data used are publicly available at PSP FIELDS (fields.ssl.berkeley.edu) and SWEAP (sweap.cfa.harvard.edu) instrument data repository.

### v2

@@ Abstract
The nature of turbulence at sub-electron scales has remained an open question, central to understanding how electrons are heated in the solar wind. This is primarily because spacecraft measurements have been limited to magnetic field fluctuations alone. We resolve this by deriving new high-resolution density fluctuations from spacecraft potential measurements of Parker Solar Probe resolving scales smaller than the electron gyro-radius ($\rho_{e}$). A systematic comparison of the density and magnetic spectra shows that both steepen near the electron scales. Notably, the density spectrum exhibits slopes close to 10/3-10/3, while the magnetic spectrum becomes consistently steeper than the density spectrum at scales smaller than $\rho_{e}$, indicating that the turbulence becomes electrostatic. These results are consistent with theoretical predictions of an electron entropy cascade, which may explain the irreversible dissipation of turbulent energy at sub-$\rho_{e}$ scales. The magnetic spectrum, however, is not as steep as expected for the electron entropy cascade, which may be due to limited signal-to-noise ratio and the presence of weakly damped electromagnetic fluctuations near $\rho_{e}$.
@@ Introduction:
Turbulence in the solar wind drives a nonlinear cascade of energy injected at large scales to smaller kinetic scales, where it gets dissipated [25, 56, 8, 12]. This is well established from observations of magnetic power spectra, which exhibit a broadband wavenumber power-law scaling $P(k)\sim k^{\alpha}$ with $\alpha\simeq-3/2$ to 5/3-5/3 at magneto-hydrodynamic (MHD) scales [15, 57, 52, 40], consistent with models of Alfvénic turbulence [21, 50, 47, 1]. At sub-ion scales, between the ion gyro-radius ($\rho_{i}$) and electron inertial scale ($d_{e}$), density ($n$) and magnetic field ($B$) spectra have been extensively measured, showing slopes close to 7/3-7/3 or 8/3-8/3 [16, 44], in agreement with predictions of kinetic Alfvén (KA) turbulence [29, 22, 10, 9, 28, 49, 23]. In the range between $d_{e}$ and $\rho_{e}$, an increase in magnetic compressibility was found, with the magnetic spectrum showing a slope near 11/3-11/3, interpreted as a transition from KA to inertial-KA cascade under low electron beta ($\beta_{e}$) conditions, in the magnetosheath [17]. In contrast, instrumental limitations have prevented a similar extent of progress at sub-electron scales (below the electron gyro-radius, $\rho_{e}$), with measurements affected by low signal-to-noise ratio (SNR) and restricted to magnetic fluctuations alone. As a result, the nature of turbulence at these scales has remained an open question.
@@
We use high-resolution search coil magnetometer (SCM) and antenna potential burst mode data from the FIELDS instrument suite on PSP, sampled at $\sim 19\,\mathrm{kHz}$ [6, 37, 31]. Data from encounters E10–E18 were surveyed for intervals with enhanced fluctuation amplitudes near the Sun, where density and magnetic spectra could be resolved at scales below $\rho_{e}$. Although $\rho_{e}$ and $d_{e}$ shift beyond instrumental resolution near perihelion, several $27\,\mathrm{s}$ intervals within $0.15–0.2\,\textrm{au}$ were identified where sub-$\rho_{e}$ scales could be resolved. We selected 76 such intervals without dust hits or wave activity, exhibiting $\textrm{SNR}\gtrsim 10$ at the observed spectral break around $d_{e}$ (detailed in the next section). The intervals used do not introduce any systematic biases and constitute a representative sample spanning a range of solar wind conditions near the Sun, with plasma beta $\beta_{i}\simeq\beta_{e}\simeq 0.4$ and a temperature ratio $T_{i}/T_{e}\simeq 1$.
Since SPAN, SPC [32], and QTN [39] density measurements have a low sampling rate ($\sim 0.14\,\mathrm{Hz}$), we use spacecraft potential ($V_{\textrm{sc}}$) as a proxy for density to resolve scales smaller than $\rho_{e}$. This method is based on the following principle: Solar illumination charges the spacecraft body positive via photoemission, attracting nearby electrons and thereby making $V_{\mathrm{sc}}$ sensitive to local electron density ($n$) [43, 34, 20, 35, 41]. The resulting balance between photoelectric and thermal electron currents leads to an exponential relation: $n\propto e^{-aV_{\mathrm{sc}}/V_{\mathrm{pe}}+C_{\mathrm{b}}}$, where $V_{\mathrm{pe}}$ is the photoelectron thermal energy. $V_{\mathrm{sc}}$ is computed as the negative average of the four PSP antenna potentials, $V_{\mathrm{sc}}=-(V_{1}+V_{2}+V_{3}+V_{4})/4$, which includes the antenna bias voltage. To account for this bias, an offset parameter $(C_{\mathrm{b}})$ was included. Other currents, such as ion or secondary electron currents, may be present but are negligible. This relation is indeed observed between electron QTN density measurements and $V_{\mathrm{sc}}$ for the intervals used. Typically at sub-ion scales, $\delta V_{\textrm{sc}}/V_{\mathrm{pe}}\ll 1$, implying $\delta n\propto\delta V_{\mathrm{sc}}$, up to first order. Thus, the two spectra share the same shape and the $V_{\textrm{sc}}$ spectrum can be used to measure the density spectral slopes.
@@
Figure 1: Power spectral density of (a) $V_{\mathrm{sc}}$ and (b) $B$ (MAG - dashed, SCM - solid). The least square fits (vertically shifted for clarity) and the break frequency ($f_{b}$) using Eq. 1 are shown. Thin-dashed lines indicate the flat analog-to-digital converter and the frequency dependent SCM noise floors. (c) $V_{\mathrm{sc}}$ and $B$ spectra compensated by the sub-ion range scaling $(f^{-\alpha_{1}}P(f))$. The compensated spectra are shown as dotted lines for SNR $\lesssim 2$.
[…]
$P(f)=\cases{A}_{1}f^{\alpha_{1}},&f<f_{b}\\ A_{2}f^{\alpha_{2}},&f\geq f_{b}$ (1)
@@
Figure 2: (a) $V_{\mathrm{sc}}$ and (b) $B$ spectra for all intervals smoothed using a variable width running mean window. Average slopes at sub-ion and sub-electron scales are shown. Blue diamonds mark the breaks obtained from fits. Mean $f_{d_{e}}$ and its variation across the intervals are shown by the red marker.
As depicted in Fig. 1, the $n$ and $B$ spectra show similar slopes of 2.77-2.77 and 2.84-2.84, respectively, in the sub-ion range ($f<f_{d_{e}}$). A spectral break is identified from the fit at frequency $f_{b}$ approximately near $f_{d_{e}}$ or $f_{\rho_{e}}$ in both spectra. Thereafter, the $B$ spectrum steepens to 4.40-4.40 while the $n$ spectrum is comparatively shallower at 3.39-3.39. This transition is clear from the compensated spectra, which steepen compared to the sub-ion range scaling ($f^{\alpha_{1}}$, horizontal dashed line) at $f_{d_{e}}$ or $f_{\rho_{e}}$. A similar trend was observed across all studied intervals as evident from Fig. 2. This figure also marks the spectral breaks (blue diamonds), showing they appear approximately around $f_{d_{e}}$. While a steepening in the $B$ spectrum has been reported before close to the electron scales, we show here that the $n$ spectrum also steepens, although to a lesser degree than the $B$ spectrum. This difference is evident from the distribution of $n$ (Fig. 3(a)) and $B$ (Fig. 3(b)) spectral slopes obtained from the fits (Eq. 1) across all intervals. In the sub-ion range ($f<f_{b}$), the distributions peak near 8/3-8/3, with mean slopes of $2.59\pm 0.02$ for $B$ and $2.55\pm 0.03$ for $n$, with the uncertainties indicating the standard error of the mean. This is consistent with previous observations [16, 44], and the theoretical and numerical predictions of KA turbulence, which suggest a slope close to 7/3-7/3 or 8/3-8/3 for both fields [29, 22, 10, 9, 28, 49, 23]. In the sub-electron range ($f>f_{b}$), both spectra steepen, but differ in slope. While $n$ shows a mean slope of $3.20\pm 0.03$, $B$ steepens to $3.78\pm 0.04$, occasionally reaching up to 4.8-4.8. This is the main result of this Letter, with this difference in slope indicating a breakdown of KA turbulence close to $\rho_{e}$.
@@ Conclusion:
We have presented high-resolution density spectra obtained from spacecraft potential measurements, alongside magnetic spectra resolving scales smaller than $\rho_{e}$, leading to the following main findings: At sub-$\rho_{e}$ scales, the density and magnetic spectra diverge, with the density spectrum exhibiting slopes close to 10/3-10/3 and the magnetic spectrum comparatively steeper. These results indicate that turbulence at these scales becomes predominantly electrostatic and are consistent with the theory of electron entropy cascade [48, 49], a phase-space process, which can explain how electrons are irreversibly heated in a wide range of space and astrophysical systems. In this context, our study motivates the need for improved particle instruments to directly measure the associated small-scale velocity-space structure and to fully understand the role of nonlinear phase-space dynamics in the irreversible dissipation of turbulent energy in weakly collisional plasmas. We also note that, while the magnetic spectrum is steeper than the density spectrum, it appears shallower than predicted by the entropy cascade theory. This discrepancy may arise from instrumental noise, allowing the possibility that the true magnetic spectrum may satisfy a steeper power law than observed. Alternatively, the shallower magnetic spectrum may result from residual, weakly damped electromagnetic fluctuations near $\rho_{e}$. Future theoretical developments and improved measurements may help resolve this discrepancy, with our results providing important quantitative constraints.
@@ Acknowledgements.
S.M. is supported by a QMUL PhD studentship. C.H.K.C. is supported by UKRI Future Leaders Fellowship MR/W007657/1. C.H.K.C. and D.M. are supported by STFC Consolidated Grant ST/X000974/1. S.M. acknowledges A.A. Schekochihin, P.A. Simon, C. Sishtla, S. Greess and F. Koller for useful discussions. The authors acknowledge helpful discussions with D. Malaspina, M. Pulupa, M. Liu and J. Bonell regarding the spacecraft potential data. Data used are publicly available at the PSP FIELDS (fields.ssl.berkeley.edu) and SWEAP (sweap.cfa.harvard.edu) data repositories.